THE SPECTROSCOPIC STUDY OF ESTROGEN AND ITS HYDRATED CLUSTERS IN A SUPER SONIC JET

\underline{FUMIYA MORISHIMA}, YOSHIYA INOKUCHI and TAKAYUKI EBATA, Department of Chemistry, Graduate School of Science, Hiroshima University, Higashi-hiroshima, 739-8526, Japan.

Structures of estrogen and its hydrated clusters have been studied by several laser spectroscopies in supersonic jet. The electronic spectrum of estrogen shows several origin bands. By observing UV-UV hole-burning and IR-UV spectra, it is concluded they are due to different conformers originating from difference of orientation of OH group(s). We also observed electronic and IR spectra of estrogen-H_{2}O. By aids of DFT calculations, the conformations and hydrated structures are determined.